OpenStudy (anonymous):

find the fourth degree polynomial with real coefficients having zeros: 0,1,2

hero (hero):

If the polynomial has real zeroes 0,1,2 then x = 0 x = 1 x = 2 Placing all terms on the left side you have x - 0 = 0 x - 1 = 0 x - 2 = 0 By zero product property (x - 0)(x - 1)(x - 2) = 0 Since we need a fourth degree polynomial, add another factor: (x - 0)(x - 1)(x - 1)(x - 2) = 0 x(x - 1)^2(x - 2) = 0

hero (hero):

x(x - 1)^2(x - 2) = f(x) Multiply that out to get your fourth degree polynomial.

OpenStudy (anonymous):

thank you so much!

hero (hero):

You're welcome.